# WetCat: Automating Skill Assessment in Wetlab Cataract Surgery Videos

### v1

@@ Abstract.
To meet the growing demand for systematic surgical training, wetlab environments have become indispensable platforms for hands-on practice in ophthalmology. Yet, traditional wetlab training depends heavily on manual performance evaluations, which are labor-intensive, time-consuming, and often subject to variability. Recent advances in computer vision offer promising avenues for automated skill assessment, enhancing both the efficiency and objectivity of surgical education. Despite notable progress in ophthalmic surgical datasets, existing resources predominantly focus on real surgeries or isolated tasks, falling short of supporting comprehensive skill evaluation in controlled wetlab settings. To address these limitations, we introduce WetCat, the first dataset of wetlab cataract surgery videos specifically curated for automated skill assessment. WetCat comprises high-resolution recordings of surgeries performed by trainees on artificial eyes, featuring comprehensive phase annotations and semantic segmentations of key anatomical structures. These annotations are meticulously designed to facilitate skill assessment during the critical capsulorhexis and phacoemulsification phases, adhering to standardized surgical skill assessment frameworks. By focusing on these essential phases, WetCat enables the development of interpretable, AI-driven evaluation tools aligned with established clinical metrics. This dataset lays a strong foundation for advancing objective, scalable surgical education and sets a new benchmark for automated workflow analysis and skill assessment in ophthalmology training. The dataset and annotations are publicly available in Synapse (https://www.synapse.org/Synapse:syn66401174/files/).
Surgical Skill Assessment, Cataract Surgery, Surgical Phase Recognition, Semantic Segmentation, Surgical Workflow Analysis, Computer-Assisted Interventions
journalyear: copyright: none
@@ 1. Introduction
Cataract surgery is one of the most commonly performed surgical procedures worldwide, with approximately four million cases annually in the United States and around 20 million globally (mcdermott2022overview, ; rossi2021cataract, ). Driven by an aging global population, the demand for cataract surgery is expected to rise significantly, with an anticipated increase of up to 128% by 2036. Concurrently, the World Health Organization projects that cataract-related blindness will affect nearly 40 million individuals by 2025 (VB2020, ). As a result, mastery of cataract surgery has become a critical component of ophthalmology training programs, which increasingly emphasize the need for systematic evaluation and improvement of surgical skills among trainees.
@@
Advances in surgical education technologies have led to the integration of simulation-based training within ophthalmology curricula. Wetlab environments, in particular, have emerged as essential platforms for hands-on surgical practice, offering anatomically realistic models without the risks associated with patient care. Unlike drylab simulations, which often incorporate embedded sensors to facilitate objective performance evaluation, wetlab training primarily relies on manual assessment by expert reviewers, a process that is labor-intensive, costly, and frequently lacking in precision and detailed feedback. In this context, automated skill assessment through computer vision presents a promising solution, offering scalable and objective evaluation methods that can enhance both training efficiency and educational outcomes.
@@
Cataract surgery itself is a technically demanding procedure that necessitates precise motor control, nuanced intraoperative decision-making, and delicate manipulation within a constrained anatomical space. Despite the critical role of wetlab training in developing surgical competencies, current publicly available datasets are derived from real patient surgeries and focus on isolated tasks such as instrument detection (al2019cataracts, ), phase recognition (LocalPhase, ), or segmentation of anatomical structures and instruments (CaDIS, ; ReCal-Net, ; ghamsarian2024deeppyramid+, ; ghamsarian2022deeppyramid, ). Other datasets target specific objectives, including irregularity detection (LensID, ; ghamsarian2023predicting, ) and relevance-based video compression (ghamsarian2020relevance, ). While these resources have advanced computer vision applications in ophthalmic surgery, they are not designed for comprehensive, automated skill assessment, particularly under the controlled conditions of wetlab environments. Furthermore, datasets based on real surgeries present significant domain adaptation challenges when applied to wetlab settings.
@@
Recognizing these gaps, we introduce WetCat, the first dataset of wetlab cataract surgery videos specifically curated for automated skill assessment. WetCat consists of high-resolution recordings of cataract procedures performed on artificial eyes by trainee ophthalmologists. Each video is meticulously annotated with surgical phase labels and semantic segmentations of key anatomical structures, following standardized assessment frameworks such as GRASIS (cremers2005global, ) and OSCAR (golnik2013development, ). The dataset focuses on the critical phases of capsulorhexis and phacoemulsification, which are essential to cataract surgery training. The standardized, reproducible nature of wetlab procedures further enhances WetCat’s utility for benchmarking and developing AI-driven evaluation tools.
In addition to supporting conventional video analysis tasks such as phase recognition and instrument tracking, WetCat enables objective and interpretable assessments of surgical proficiency. By aligning with established clinical skill metrics, it paves the way for intelligent, data-rich training systems aimed at improving the quality, consistency, and scalability of ophthalmic surgical education.
@@
The remainder of this paper is organized as follows. Section 2 reviews the skill assessment criteria that motivate the need for our phase annotations and surgical scene segmentations. Section 3 describes the WetCat dataset in detail. Section 4 presents experimental validations of the annotations. Finally, Section 5 concludes the paper.
@@
Figure 1. Sample frames from relevant phases in a wetlab cataract surgery.
@@
Sample frames from relevant phases in a wetlab cataract surgery.
@@ 2. Methods
In this section, we present the evaluation metrics for two critical phases of cataract surgery: capsulorhexis and phacoemulsification. These skill metrics are derived from established standards such as GRASIS (cremers2005global, ) and OSCAR (golnik2013development, ), ensuring clinical relevance and standardization. The dataset annotations are structured in alignment with these metrics to facilitate objective skill assessment. Figure 1 illustrates the overall framework for skill assessment, highlighting the role of phase recognition and semantic segmentation in evaluating surgical skill during the two targeted phases.
@@ 2.1. Capsulorhexis Assessment Metrics
  • Roundness: Measured by the circularity ratio, defined as
    $\frac{4\pi\cdot\text{Area}}{\text{Perimeter}^{2}},$
    where values closer to 1 indicate a near-perfect circle.
  • Centration: Assesses alignment accuracy by computing the Euclidean distance between the centers of the capsulorhexis and the iris, normalized by the iris radius:
    Centration $\text{Centration Metric}=\frac{\|\text{Capsulorhexis Center}-\text{Iris Center}\|}{\text{Iris Radius}}.$
@@ 2.2. Phacoemulsification Assessment Metrics
  • Eye Stability: Assessed by tracking the iris center throughout the phase. Deviations from the initial position are expressed as a percentage of the iris diameter, with deviations within ±$\pm 10\%$ considered acceptable. Larger displacements may indicate inadequate fixation or excessive instrument force.
@@ 3. Dataset
The WetCat dataset comprises 60 cataract surgery videos recorded between 2024 and 2025, performed by junior to mid-level surgeons using the Haag-Streit OSTC microscope system. Each surgery was conducted under a binocular surgical microscope, providing a magnified, illuminated, and stereoscopic view of the operative field. Surgeons adjusted the microscope’s focus to maintain optimal clarity, while an integrated camera system captured high-resolution video recordings of the complete procedures for comprehensive post-operative analysis and skill evaluation.
Each video in the dataset captures the complete surgical workflow, with an average duration of 794 seconds and a standard deviation of 509 seconds. In addition to the full video recordings, the dataset is comprehensively annotated with surgical phase labels for all cases. Specifically, each video is segmented into four distinct phases: (1) Capsulorhexis (referred to as Rhexis), (2) Phacoemulsification (Phaco), (3) Idle, and (4) Rest. Figure 2 summarizes the distribution of phases across all videos and reports the overall percentage representation of each phase within the dataset. Furthermore, Table 2 presents normalized phase annotations for a subset of 20 representative videos, providing a detailed overview of phase transitions and durations.
@@
This comprehensive annotation framework enables detailed analysis of both surgical workflow and scene composition, supporting a wide range of computer vision tasks relevant to automated surgical skill assessment. Table 1 provides a comparative summary of the WetCat dataset alongside existing datasets focused on cataract surgery phase recognition and semantic segmentation.
@@
Attribute Type CaDIS (CaDIS, ) CatRel (LocalPhase, ) Cataract-1K (Cataract-1K_dataset, ) WetCat
Acquisition Period 2015 2017–2018 2021–2023 2024–2025
Phase Annotations # Videos 22 56 60
Segment Annotations # Frames 4,670 2,256 1,469
Phase Classes 5 13 4
Segment Classes Anatomy N/A
Instruments N/A
Rhexis N/A
Frame Rate (fps) N/A 25 30 30
Table 2. Visualization of relevant phase annotations for ten representative wetlab videos from our dataset.
@@
Figure 4. Sample frames from relevant phases in a wetlab cataract surgery.
@@
Sample frames from relevant phases in a wetlab cataract surgery.
@@
The datasets are licensed under CC BY-NC-ND. We provide all code for mask creation using XML annotations and phase extraction using CSV files, as well as the training IDs for four-fold validation and usage instructions in the GitHub repository of the paper (https://github.com/Negin-Ghamsarian/WetCat).
@@ 4. Technical Validation
In this section, we validate the quality and consistency of our multi-task annotations by training a set of state-of-the-art neural network architectures, each tailored to its respective task. The performance of these models is thoroughly assessed using task-appropriate evaluation metrics, providing a robust measure of the accuracy and reliability of our annotated data.
@@ 4.1. Experimental Settings for Phase Recognition
We evaluate three types of neural network architectures for phase recognition. First, we adopt a combined CNN-RNN framework, where the CNN backbone is VGG16, and the RNN component is configured with four variants: GRU, LSTM, BiGRU, and BiLSTM, following the architectures proposed in (LocalPhase, ). Second, we implement a vision transformer model that integrates VGG16 as the backbone with a lightweight transformer head comprising 4.30 M parameters, as described in (nasirihaghighi2024event, ). Third, we employ a 3D convolutional residual network (ResNet3D-18) following (ResNet3D, ). All backbone networks are pre-trained on ImageNet (ImageNet, ), except for ResNet3D, which is trained from scratch.
For phase recognition, we merge the ”Idle” and ”Rest” phases into a single class, resulting in a three-class segmentation task. Phase recognition is conducted on three-second video clips by randomly sampling 10 frames from 90 frames within each clip to construct the input sequences, as per the protocols in (LocalPhase, ; 10178763, ). We evaluate the performance using accuracy, macro recall, and macro F1 score as the primary metrics.
@@
Adapt-Net (LensID, ) VGG16 24.69 M Surgical Videos
UNet+++++ + (UNet++, ) VGG16 24.24 M Medical Images
ReCal-Net (ReCal-Net, ) VGG16 22.93 M Surgical Videos
CPFNet (CPFNet, ) VGG16 39.17 M Medical Images
CE-Net (CE-Net, ) ResNet34 29.90 M Medical Images
DeepLabV3+ (DeepLabV3, ) ResNet50 26.68 M Natural Images
UPerNet (UPerNet, ) ResNet50 51.26 M Natural Images
U-Net+ (U-Net, )  [^1] Note that UNet+ is an improved version of UNet, where we use VGG16 as the backbone network and double convolutional blocks (two consecutive convolutions followed by batch normalization and ReLU layers) as decoder modules. VGG16 22.55 M Medical Images
SAM (kirillov2023segment, ) ViT-B 91 M Natural Images
@@ 4.2. Experimental Settings for Semantic Segmentation
We conduct a series of experiments to assess the reliability of our pixel-level annotations, benchmarking against multiple state-of-the-art methods designed for general imagery, medical imaging, and surgical video analysis. Detailed descriptions of these baseline approaches are provided in Table 3.
Given the substantial number of parameters in the Segment Anything Model (SAM) and the considerable computational resources required for full training, we adopt an efficient fine-tuning strategy. Specifically, we freeze both the vision encoder and the prompt encoder, and investigate two training configurations: (1) training only the mask decoder, resulting in 4,058,340 trainable parameters, and (2) applying Low-Rank Adaptation (LoRA) (hu2022lora, ) to the linear and convolutional (Conv2D) layers of the vision encoder, with a rank of 16 ($r=16$), scaling factor $\alpha=32$, and a dropout rate of 0.1, increasing the trainable parameters to 6,647,012. Unless otherwise specified, SAM is fine-tuned using grid prompts as input.

### v3

# WetCat: Enabling Automated Skill Assessment in Wet-Lab Cataract Surgery Videos

@@ Abstract.
To meet the growing demand for systematic surgical training, wet-lab environments have become indispensable platforms for hands-on practice in ophthalmology. Yet, traditional wet-lab training depends heavily on manual performance evaluations, which are labor-intensive, time-consuming, and often subject to variability. Recent advances in computer vision offer promising avenues for automated skill assessment, enhancing both the efficiency and objectivity of surgical education. Despite notable progress in ophthalmic surgical datasets, existing resources predominantly focus on real surgeries or isolated tasks, falling short of supporting comprehensive skill evaluation in controlled wet-lab settings. To address these limitations, we introduce WetCat, the first dataset of wet-lab cataract surgery videos specifically curated for automated skill assessment. WetCat comprises high-resolution recordings of surgeries performed by trainees on artificial eyes, featuring comprehensive phase annotations and semantic segmentations of key anatomical structures. These annotations are meticulously designed to facilitate skill assessment during the critical capsulorhexis and phacoemulsification phases, adhering to standardized surgical skill assessment frameworks. By focusing on these essential phases, WetCat enables the development of interpretable, AI-driven evaluation tools aligned with established clinical metrics. This dataset lays a strong foundation for advancing objective, scalable surgical education and sets a new benchmark for automated workflow analysis and skill assessment in ophthalmology training. The dataset and annotations are publicly available in Synapse (https://www.synapse.org/Synapse:syn66401174/files/).
Surgical Skill Assessment, Cataract Surgery, Surgical Phase Recognition, Semantic Segmentation, Surgical Workflow Analysis, Computer-Assisted Interventions, Wet-Lab Cataract Surgery, Cataract Surgery Dataset.
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758274isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Image and video acquisitionccs: Computing methodologies Activity recognition and understandingccs: Computing methodologies Image segmentation
Figure 1. Hands-on cataract surgery training in the wet lab.
[…]
Hands-on cataract surgery training in the wet lab.
@@ 1. Introduction
Cataract surgery is one of the most commonly performed surgical procedures worldwide, with approximately four million cases annually in the United States and around 20 million globally (McDermott and Liang, 2022; Rossi et al., 2021). Driven by an aging global population, the demand for cataract surgery is expected to rise significantly, with an anticipated increase of up to 128% by 2036. Concurrently, the World Health Organization projects that cataract-related blindness will affect nearly 40 million individuals by 2025 (Burton et al., 2021). As a result, mastery of cataract surgery has become a critical component of ophthalmology training programs, which increasingly emphasize the need for systematic evaluation and improvement of surgical skills among trainees.
@@
Advances in surgical education technologies have led to the integration of simulation-based training within ophthalmology curricula. Wet-lab environments, in particular, have emerged as essential platforms for hands-on surgical practice, offering anatomically realistic models without the risks associated with patient care (Figure 1). Unlike dry-lab simulations, which often incorporate embedded sensors to facilitate objective performance evaluation, wet-lab training primarily relies on manual assessment by expert reviewers, a process that is labor-intensive, costly, and frequently lacking in precision and detailed feedback. In this context, automated skill assessment through computer vision presents a promising solution, offering scalable and objective evaluation methods that can enhance both training efficiency and educational outcomes.
@@
Despite the critical role of wet-lab training in developing surgical competencies, current publicly available datasets are derived from real patient surgeries (Ghamsarian et al., 2024a) and in most cases focus on isolated tasks such as instrument detection (Al Hajj et al., 2019), phase recognition (Ghamsarian et al., 2021a), or segmentation of anatomical structures and instruments (Grammatikopoulou et al., 2021). Other datasets target specific objectives, including irregularity detection (Ghamsarian et al., 2021b, 2024b) and relevance-based video compression (Ghamsarian et al., 2020a). Accordingly, a considerable body of research has focused on content analysis in real-world cataract surgery (Nasirihaghighi et al., 2025; Ghamsarian et al., 2024c, 2022; Marafioti et al., 2021; Hu et al., 2024b; Ghamsarian et al., 2025; Nasirihaghighi, 2025; Ghamsarian et al., 2020b; Ghamsarian, 2020; Hu et al., 2024a). While these resources have advanced computer vision applications in ophthalmic surgery, they are not designed for comprehensive, automated skill assessment, particularly under the controlled conditions of wet-lab environments. Furthermore, datasets based on real surgeries present significant domain adaptation challenges when applied to wet-lab settings (Ghamsarian et al., 2023).
@@
Recognizing these gaps, we introduce WetCat, the first dataset of wet-lab cataract surgery videos specifically curated for automated skill assessment. WetCat consists of high-resolution recordings of cataract procedures performed on artificial eyes by trainee ophthalmologists. Each video is meticulously annotated with surgical phase labels and semantic segmentations of key anatomical structures, following standardized assessment frameworks such as GRASIS (Cremers et al., 2005) and OSCAR (Golnik et al., 2013). The dataset focuses on the critical phases of capsulorhexis and phacoemulsification, which are essential to cataract surgery training. The standardized, reproducible nature of wet-lab procedures further enhances WetCat’s utility for benchmarking and developing AI-driven evaluation tools. In addition to supporting conventional video analysis tasks such as phase recognition and instrument tracking, WetCat enables objective and interpretable assessments of surgical proficiency. By aligning with established clinical skill metrics, it paves the way for intelligent, data-rich training systems aimed at improving the quality, consistency, and scalability of ophthalmic surgical education.
@@
The remainder of this paper is organized as follows. Section 2 reviews the skill assessment criteria that motivate the need for our phase annotations and surgical scene segmentations. Section 3 describes the WetCat dataset in detail. Section 4 presents benchmarking and experimental validations of the annotations. Finally, Section 5 summarizes the work and concludes the paper.
@@
Figure 2. Sample frames from relevant phases in a wet-lab cataract surgery.
@@
Sample frames from relevant phases in a wet-lab cataract surgery.
@@ 2. Methods
In this section, we present the evaluation metrics for two critical phases of cataract surgery: capsulorhexis and phacoemulsification. These skill metrics are derived from established standards, including GRASIS (Cremers et al., 2005) and OSCAR (Golnik et al., 2013), ensuring clinical relevance and standardization. The dataset annotations are structured in alignment with these metrics to facilitate objective skill assessment. Figure 2 illustrates the overall framework for skill assessment, highlighting the role of phase recognition and semantic segmentation in evaluating surgical skill during the two targeted phases.
@@ 2.1. Capsulorhexis Assessment Metrics
  • Roundness: Measured by the circularity ratio, which compares the object’s area to the square of its perimeter, with values near 1 indicating a nearly perfect circle.
  • Centration: Assesses alignment accuracy using a centration metric, defined as the Euclidean distance between the capsulorhexis and limbus centers normalized by the limbus radius.
@@ 2.2. Phacoemulsification Assessment Metrics
  • Eye Stability: Assessed by tracking the limbus center throughout the phase. Deviations from the initial position are expressed as a percentage of the limbus diameter, with deviations within $\pm 10\%$ considered acceptable. Larger displacements may indicate inadequate fixation or excessive instrument force.
@@ 3. Dataset
The WetCat dataset comprises 60 cataract surgery videos recorded during 2024–2025, performed by junior to mid-level surgeons using the Haag-Streit OSTC microscope system. Each surgery was conducted under a binocular surgical microscope, providing a magnified, illuminated, and stereoscopic view of the operative field. Surgeons adjusted the microscope’s focus to maintain optimal clarity, while an integrated camera system captured high-resolution video recordings of the complete procedures for comprehensive post-operative analysis and skill evaluation.
Each video in the dataset captures the complete surgical workflow, with an average duration of 794 seconds and a standard deviation of 509 seconds. In addition to the full video recordings, the dataset is comprehensively annotated with surgical phase labels for all cases. Specifically, each video is segmented into four distinct phases: (1) Capsulorhexis (referred to as Rhexis), (2) Phacoemulsification (Phaco.), (3) Idle, and (4) Rest. Figure 3 summarizes the distribution of phases across all videos and reports the overall percentage of each phase within the dataset. Furthermore, Table 1 presents normalized phase annotations for a subset of 20 representative videos, providing a detailed overview of phase transitions and relative durations.
@@
This comprehensive annotation framework enables detailed analysis of both surgical workflow and scene composition, supporting a wide range of computer vision tasks relevant to automated surgical skill assessment. Table 2 provides a comparative summary of the WetCat dataset alongside existing datasets focused on phase recognition and semantic segmentation in cataract surgery.
@@
Table 1. Visualization of relevant phase annotations for ten representative wet-lab videos from our dataset.
@@
Figure 5. Sample frames from relevant phases in a wet-lab cataract surgery.
@@
Sample frames from relevant phases in a wet-lab cataract surgery.
Attribute Type CaDIS (Grammatikopoulou et al., 2021) CatRel (Ghamsarian et al., 2021a) cataract-1k (Ghamsarian et al., 2024a) WetCat
Acquisition Period 2015 2017–2018 2021–2023 2024–2025
Phase Annotations # Videos 22 56 60
Segment Annotations # Frames 4,670 2,256 1,469
Phase Classes 5 13 4
Segment Classes Anatomy N/A
Instruments N/A
Rhexis N/A
Frame Rate (fps) N/A 25 30 30
@@
The datasets are licensed under CC BY-NC-SA 4.0. We provide all code for annotation processing, as well as the training IDs for four-fold validation and usage instructions in the GitHub repository of the paper (https://github.com/Negin-Ghamsarian/WetCat).
@@ 4. Technical Validation
In this section, we validate the quality and consistency of our multi-task annotations by training a set of state-of-the-art neural network architectures, each tailored to its respective task. The performance of these models is thoroughly assessed using task-appropriate evaluation metrics, providing a robust measure of the accuracy and reliability of our annotations.
@@ 4.1. Experimental Settings for Phase Recognition
We evaluate three types of neural network architectures for phase recognition. First, we adopt a combined CNN-RNN framework, where the CNN backbone is VGG16, and the RNN component is configured with four variants: GRU, LSTM, BiGRU, and BiLSTM, following the architectures proposed in (Ghamsarian et al., 2021a). Second, we implement a vision transformer model that integrates VGG16 as the backbone with a lightweight transformer head comprising 4.30 M parameters, as described in (Nasirihaghighi et al., 2024). Third, we employ a 3D convolutional residual network (ResNet3D-18) following (Tran et al., 2015). All backbone networks are pre-trained on ImageNet (Deng et al., 2009), except for ResNet3D, which is trained from scratch.
For phase recognition, we merge the ”Idle” and ”Rest” phases into a single class, resulting in a three-class segmentation task. Phase recognition is conducted on three-second video clips by randomly sampling 10 frames from 90 frames within each clip to construct the input sequences, as per the protocols in (Ghamsarian et al., 2021a; Nasirihaghighi et al., 2023). We evaluate the performance using accuracy, macro recall, and macro F1 score as the primary metrics.
@@
Adapt-Net (Ghamsarian et al., 2021b) VGG16 24.69 M Surgical Videos
UNet++++ (Zhou et al., 2020) VGG16 24.24 M Medical Images
ReCal-Net (Ghamsarian et al., 2021c) VGG16 22.93 M Surgical Videos
CPFNet (Feng et al., 2020) VGG16 39.17 M Medical Images
CE-Net (Gu et al., 2019) ResNet34 29.90 M Medical Images
DeepLabV3+ (Chen et al., 2018) ResNet50 26.68 M Natural Images
UPerNet (Xiao et al., 2018) ResNet50 51.26 M Natural Images
U-Net+ (Ronneberger et al., 2015)  [^1] Note that UNet+ is an improved version of UNet, where we use VGG16 as the backbone network and double convolutional blocks (two consecutive convolutions followed by batch normalization and ReLU layers) as decoder modules. VGG16 22.55 M Medical Images
SAM (Kirillov et al., 2023) ViT-B $\approx 91$ M Natural Images
@@ 4.2. Experimental Settings for Semantic Segmentation
We conduct a series of experiments to assess the reliability of our pixel-level annotations, benchmarking against multiple state-of-the-art methods designed for natural images, medical images, and surgical videos. Detailed descriptions of these baseline approaches are provided in Table 3.
Given the substantial number of parameters in the Segment Anything Model (SAM) and the considerable computational resources required for full training, we adopt an efficient fine-tuning strategy. Specifically, we freeze both the vision encoder and the prompt encoder, and investigate two training configurations: (1) training only the mask decoder, resulting in 4,058,340 trainable parameters, and (2) applying Low-Rank Adaptation (LoRA) (Hu et al., 2022) to the linear and convolutional (Conv2D) layers of the vision encoder, with a rank of 16 ($r=16$), scaling factor $\alpha=32$, and a dropout rate of 0.1, increasing the trainable parameters to 6,647,012. Unless otherwise specified, SAM is fine-tuned using grid prompts as input.